OpenStudy (anonymous):

Using elimination method solve this: a + 6b = 10 a + 2b= -2 Help please! Struggle with these!

OpenStudy (abb0t):

Multiply the 2nd equation by -1 and eliminate the 'a' term to get: -3b = 12 Now you have the value for b. Next, solve for a by using the value you got for "b"

OpenStudy (anonymous):

ok so b= -4 Then, a + 2(-4)= -2 -8= -2 a= 4? i apologize if I completely messed this up, but I wasn't kidding when i said I struggle with this! meaning math in general!!

OpenStudy (ghazi):

just put a= -2-2b in the equation a+6b=10 (-2-2b)+6b=10 -2-2b+6b=10 -2+4b=10 4b=12 b=12/ 3 and then you can put value of b in equation a=-2-2b to find a ;)

OpenStudy (anonymous):

ok so a= -2-2(4) a=-2-8 a=10

OpenStudy (abb0t):

Correct.